### v2

# Genetic algorithm and edge-colorings of complete graphs with connected classes

Abstract
In this study, the Rank Genetic Algorithm (Rank GA) was adapted to address a problem in the field of Chromatic Graph Theory: the connected pseudo-pseudoachromatic index. We successfully improved several previously known lower bounds of that index. To validate the effectiveness of these new bounds, we performed an analytical approximation, confirming their validity. The findings highlight the interdisciplinary impact of the use of genetic algorithms in theoretical mathematical research, particularly in the solution of complex combinatorial problems. Our work demonstrates that computational methods can solve problems that are intractable to humans, leading to scientific advances in many areas. Moreover, the problem we addressed showcases the robustness of the Rank GA in finding optimal solutions within a reasonable time frame, as it consistently avoids getting stuck in local optima.
Keywords: Genetic Algorithm, Rank GA, Chromatic Graph Theory, Edge-Coloring, Connected Classes
Mathematics Subject Classification: 05C15, 05C85
1 Introduction
Genetic Algorithms (GA) have emerged as powerful tools, taking advantage of efficient strategies to address complex optimization problems prevalent in real-world applications in diverse industries, including transportation and manufacturing [9]. Although GAs are widely recognized for their effectiveness in addressing practical engineering challenges, their interdisciplinary potential extends to the realm of theoretical advancements in mathematical research.
In our study, we explore the application of a particular GA: the Rank Genetic Algorithm (Rank GA) introduced in [3] to address a graph theory problem. The Rank GA incorporates genetic operators that facilitate both local search (exploitation) and global search (exploration) simultaneously. This unique feature allows for a delicate balance between exploration and exploitation. With dedicated individuals focusing on exploration, the algorithm can escape local optima. Simultaneously, during local searches, the genes of the best individuals are exploited to refine the overall solution, contributing to the algorithm’s effectiveness. Due to its versatility, the Rank GA has proven to be highly effective in solving combinatorial optimization problems, as demonstrated in successful applications [4, 5, 8].
On the other hand, within the field of mathematics, particularly Graph Theory, we focus on graph coloring research, exploring the specialized field of Chromatic Graph Theory. Our specific attention is directed towards the connected-pseudoachromatic index, denoted $\psi_{c}^{\prime}(K_{n})$. Determining $\psi_{c}^{\prime}(K_{n})$ is difficult, and so far, there have been very few exact results for $\psi_{c}^{\prime}(K_{n})$ of complete graphs with connected classes. However, it has been possible to bound it using analytical methods [2].
In this interdisciplinary study, our objective was to refine the known bounds of $\psi_{c}^{\prime}(K_{n})$ exploiting the capabilities of the Rank GA. We successfully improved several of the previously known lower bounds. To validate the effectiveness of these new bounds, we performed an analytical approximation, confirming their validity. This synergy between Genetic Algorithms and mathematical theory contributes to the intersection of practical problem-solving and theoretical advancements.
The structure of this paper is as follows. In Section 2, we provide an overview of previous work in which a GA has been successfully used to solve graph theory problems. In Section 3, we state the problem and include the necessary definitions to understand it. In Section 4, we describe the Rank GA and in Section 5, we present the obtained results. Finally, Section 7 contains the conclusions.
2 Background
In recent decades, GAs have emerged as effective tools to address problems of graph theory with practical applications. Specifically, various GAs have been developed to tackle the Graph Coloring Problem (GCP), a fundamental issue in graph theory. In the simplest instance of the GCP, the objective is to assign colors to the vertices of a graph in such a way that no two adjacent vertices share the same color. The goal is to find a valid coloring of the graph with the minimal number of colors. Given its NP-hard nature and numerous practical applications, the GCP stands out as the most extensively researched combinatorial optimization problem in both computer science and mathematics [7]. In [12], you can find the most significant work in which GAs have been used to address the GCP. In addition to GCP, GAs have also been applied to other graph problems, among the most recent works are: the Shortest Path Problem [6], the Traveling Salesman Problem [13], Maximum Matching in bipartite graphs [11], and the Maximum Edge-Disjoint Path Problem [9].
Although GAs have demonstrated effectiveness in the area of graph theory for solving practical problems, to the best of our knowledge, their utilization for theoretical problems is not as widespread. We are aware of the following studies where a genetic algorithm played a crucial role in addressing theoretical challenges: In [5] the Rank GA was used to find rainbow $t$-colorings of the family of Moore cages with girth 6: $(k;6)$, achieving an improvement in the upper bound for the 4-rainbow coloring of a $(4;6)$-cage. In addition, in [8], the Rank GA was adapted with recombination and mutation operators from graph theory. These operators enabled the discovery of Hajós operations to construct the symmetric 5-cycle from the complete symmetric digraph of order 3 using only Hajós operations. Finally, in [4] real solutions were identified by a genetic algorithm, thus contributing to the validation of a conjecture. These studies illustrate that GAs can make valuable contributions to the field of mathematics. This, in turn, inspired us to tackle the challenge of enhancing bounds in edge-colorings of complete graphs with connected classes.
3 Problem Statement
The maximum number of colors in a connected complete coloring of a complete graph $K_{n}$ of order $n$ that can be achieved is called the connected pseudo-achromatic index and is denoted $\psi_{c}^{\prime}(K_{n})$. There are certain values of $n$ for which it is known, and for others, only the lower and upper bounds are known [1, 2]. The problem we address here, in particular, consists in improving the known bounds [2] for some values of $n$.
3.1 Edge-Colorings of a Complete Graph
A graph is defined as a set of vertices $V$ and a set of edges $E$, and is denoted by $G=(V,E)$. It is important to note that, in a graph, there cannot be repeated edges or edges connecting a vertex to itself, since these conditions would violate the definition of a graph. A complete graph is a graph in which each pair of vertices is connected by an edge.
A complete edge-coloring in a complete graph is an assignment of colors to edges such that, for every pair of colors, there is at least a common vertex. Figure 1 (left) shows an example of a complete edge-coloring of $K_{4}$, while Figure 1 (right) shows an example of a noncomplete coloring, since the color pair (“dotted”, “dashed”) does not have a common vertex.
Figure 1: Edge-colorings in the complete graph $K_{4}$. (Left) A complete coloring with 4 colors. (Right) A noncomplete coloring with 5 colors.
A chromatic class is the set of all edges of a colored graph that have the same color. The induced subgraph of a chromatic class $X$ is formed by the edges of the chromatic class $X$ and the vertices connected to those edges.
A graph is connected when there is at least one path between each pair of its vertices. A chromatic class is connected if its induced subgraph is connected. Each chromatic class in Figure 1 is connected. A coloring is connected if all its chromatic classes are connected chromatic classes.
The connected-pseudoachromatic index $\psi^{\prime}_{c}(G)$ of a graph $G$ is the maximum number of colors $k$ for which a connected complete edge-coloring exists in the graph $G$ using $k$ colors, when $G$ is connected; otherwise, it is defined as the maximum index over its connected components. When we color vertices and each chromatic class is connected, the parameter is also known as the Hadwiger number of $G$.
When a graph is complete, the connected-pseudoachromatic index is denoted by $\psi_{c}(n)$ where $n$ is the number of vertices in the complete graph $K_{n}$.
In [2] it is shown that $\psi_{c}(n)$ can be bounded by analytical methods obtaining that $\psi_{c}(n)=\theta(n^{3/2})$. Table 1 shows the known bounds of $\psi_{c}(n)$, for complete connected graphs of $2\leq n\leq 31$ vertices. With the use of the computer, complete colorings can be achieved, and thus improve the lower bound of $\psi_{c}(n)$.
$n$ 2 3 4 5 6 7 8 9 10 11 Upper 1 3 4 6 7 10 14 18 22 25 Lower 1 3 4 6 7 10 11 12 13 14 $n$ 12 13 14 15 16 17 18 19 20 21 Upper 28 31 34 37 40 45 51 57 63 70 Lower 19 26 27 28 29 30 31 32 33 42 $n$ 22 23 24 25 26 27 28 29 30 31 Upper 74 78 82 86 90 94 98 102 106 116 Lower 43 44 45 46 47 48 49 50 51 93
Table 1: Values for $2\leq n\leq 7$ and lower bounds for $8\leq n\leq 12$ given in [1]. The other values were given in [2].
4 Description of the Rank GA
Genetic algorithms draw inspiration from the genetic evolution process observed in living organisms, addressing challenges in combinatorial optimization, network design, routing, scheduling, location and allocation, reliability design, and logistics [9]. Leveraging this evolutionary concept, we use the Rank GA to solve the problem described in Section 3, making use of its ability to escape local optima and refine solutions. The Rank GA operates by evaluating and ranking the population before applying each genetic operator. This ranking ensures that the corresponding genetic operator is uniquely applied to each individual based on their position in the ordered population. This approach enhances the ability to search for optimal solutions and to improve the bounds obtained thus far.
Individuals recombine with others of similar rank, so if one individual is fit due to possessing certain beneficial genes and another individual is fit for having different advantageous genes, their recombination can yield individuals that combine both sets of genes. If two individuals are unfit, they are likely to be quite different because there are numerous ways to be unfit. When recombined, it is highly probable that the offspring will also be distinct, exploring areas distant from the parents and, thereby, increasing the likelihood of escaping local optima. The fittest individuals tend to propagate their genes more within the population than the less fit ones. This allows the concentration of the population around the best individual, guiding the population towards the best individuals by exploiting their genes more.
In summary, with the Rank GA the worst individuals are dedicated to exploration making possible to escape from the local optima and, at the same time, the best individuals perform local search exploiting their genes to refine the best solution obtained so far. Therefore, the Rank GA is particularly suitable for optimization problems where finding a global optimum is challenging due to complex search spaces, as in the problem that we address in this study.
4.1 The Genetic Operators of the Rank GA
The Rank GA follows a specific process, starting with random initialization of the population. Subsequently, the individuals are evaluated and ranked from the best to the worst based on their fitness. The rank of the $i$-th individual in the ordered population, denoted as $r_{i}$, is determined using equation (4.1).
$\ r_{i}=\frac{i}{N-1}$ (4.1)
where $i$ ranges from 0 to $N-1$, and $N$ represents the number of individuals in the population.
Once individuals are ranked based on their fitness, from best to worst, the Rank GA applies the following operations iteratively until a stopping criterion is met:
  • Rank-based selection
  • Sort and rank
  • Rank-based recombination
  • Evaluate, sort, and rank
  • Rank-based mutation
  • Evaluate, sort, and rank
4.1.1 Rank-Based Selection
Rank-based selection involves the cloning of individuals according to a two-step procedure. The desired number of clones, cloneNbr, is calculated for each individual $i$ using Equation (4.2):
$\ \text{cloneNbr}_{i}=S\cdot(1-r_{i})^{S-1}$ (4.2)
Here, $r_{i}$ represents the rank of individual $i$, ranging from 0 to 1, and $S$ corresponds to the selective pressure of the rank-based selection operator. In the first step, the floor of $cloneNbr_{i}$ is taken, resulting in the number of clones to generate for each individual $i$. In the second step, additional clones are produced as follows: the procedure starts with $i=0$ and, while the total number of clones is less than the original number of individuals $N$, the fractional part of $cloneNbr_{i}$ is used as the probability of producing an additional individual clone $i$. If a random number in the range [0,1) is less than that probability, a clone of individual $i$ is generated. The value of $i$ is increased by modulo $N$ to ensure that it passes through all individuals.
4.1.2 Rank-Based Recombination
The Rank-Based Recombination operator forces mating between individuals with indices $i$ and $i$+1 in the ordered population, where $i$ increases by 2. This strategy ensures that individuals mate with others who are close in terms of rank. Using rank-based recombination with a selective pressure value of $S=3$, the Rank GA achieves a balance between preserving the advantageous genes of the fittest individual and introducing local search through recombination with other fit individuals.
4.1.3 Mutation by Rank
The mutation probability assigned to each individual, denoted as $p_{i}$, is determined by a monotonically increasing function of the rank. The function is defined by Equation (4.3) as follows:
$p_{i}=p_{\text{max}}\cdot r_{i}^{(\frac{\ln(p_{\text{max}}\cdot G)}{\ln(N-1)})}$ (4.3)
In this equation, $p_{\text{max}}$ represents the maximum mutation probability that can be assigned, $r_{i}$ is the rank of individual $i$, $G$ is the genotype size, and $N$ is the population size. The function assigns a mutation probability of 0 to the best individual ($r_{0}=0$), $1/G$ to the second-best individual, and $p_{\text{max}}$ to the worst individual.
5 Results
Table 2 presents the improved lower bounds obtained through the Rank GA for various values of $n$. The known upper bounds for $\psi_{c}^{\prime}(K_{n})$ are also shown for reference in Table 2. By comparing the improved lower bounds with the existing lower and upper bounds, we observed significant reductions in the gap between them.
$n$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$ Upper $1$ $3$ $4$ $6$ $7$ $10$ $14$ $18$ $22$ $25$ Lower Improved $13$ $16$ $18$ Lower $1$ $3$ $4$ $6$ $7$ $10$ $11$ $12$ $13$ $14$ $n$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$ $21$ Upper $28$ $31$ $34$ $37$ $40$ $45$ $51$ $57$ $63$ $70$ Lower Improved $21$ $32$ $35$ $37$ $39$ $43$ Lower $19$ $26$ $27$ $28$ $29$ $30$ $31$ $32$ $33$ $42$ $n$ $22$ $23$ $24$ $25$ $26$ $27$ $28$ $29$ $30$ $31$ Upper $74$ $78$ $82$ $86$ $90$ $94$ $98$ $102$ $106$ $116$ Lower Improved $46$ $50$ $52$ $54$ $55$ $59$ $62$ $64$ $66$ Lower $43$ $44$ $45$ $46$ $47$ $48$ $49$ $50$ $51$ $93$
Table 2: Improved lower bounds with the Rank GA for $\psi_{c}(n)$.
Figure 2 visually represents the values detailed in Table 2. As can be seen, the known lower bound presents sudden steps at certain graph sizes, while the improved lower bounds present smoother growth.
[…]
Figure 2: Improved lower bounds with the Rank GA for $\psi_{c}(n)$.
6 On the Upper Bound
In [2] it is shown that a complete graph of $n$ vertices accepts a coloring with $(\sqrt{n-1})^{3}/2$ colors and that a coloring with $(n-1)(\sqrt{n/2+1/16}+1/4)$ colors or more cannot be complete and connected at the same time.
On the one hand, the asymptotic growth of $\psi_{c}(n)$ is $\Theta(n^{\frac{3}{2}})$, on the other hand, the coefficient of the main term of the lower bound is $\frac{1}{2}$, while it is $\frac{1}{\sqrt{2}}$ for the upper bound. This gives a significant difference in the values in Table 1.
A natural question is: Are the bounds obtained by the Rank GA good? For this, we give an approximation of the upper bound of the $\psi_{c}(n)$ function by doing an analysis of the average degree with which two acyclic chromatic classes are incident.
Experimentally, only some chromatic classes contain cycles. Table 3 shows the average degrees of incidence between chromatic classes for $n\leq 30$. To achieve a range of heuristically acceptable values, we obtain an approximation of $\psi_{c}(n)$ under the hypothesis that each chromatic class contains at least one cycle, the average degree of which is at least 2.
$n$ num colors average degrees of incidence
8 11 1.563636
9 13 1.576923
10 16 1.391666
11 18 1.398692
12 21 1.361904
13 26 1.120000
14 27 1.333333
15 28 1.550264
16 29 1.894088
17 32 1.616935
18 35 1.591596
19 37 1.696696
20 39 1.716599
21 43 1.660022
22 46 1.662801
23 50 1.651428
24 52 1.718702
25 54 1.758909
26 55 1.806060
27 59 1.832846
28 62 1.875198
29 64 1.875000
30 66 1.920279
Table 3: The average degrees of incidence between the chromatic classes obtained by the Rank GA.
Theorem 6.1.
If $n\geq 8$, then an approximate upper bound is
$\psi_{c}(n)\leq k\approx\frac{n(n-1)}{\sqrt{4n-3}-1}=\frac{1}{2}n^{\frac{3}{2}}+O(n),$
Furthermore, under the hypothesis that each class contains at least one cycle (if $n\geq 9$), then an upper bound heuristic is
$\psi_{c}(n)\leq k\approx\frac{n^{\frac{1}{2}}(n-1)}{2}=\frac{1}{2}n^{\frac{3}{2}}-O(n^{\frac{1}{2}}).$
Proof.
Let $n\geq 8$ and $c\colon E(K_{n})\rightarrow\{1,...,k\}$ a connected and complete edge-coloring of $K_{n}$ with $k=\psi_{c}(n)$ colors.
Let $x$ be the cardinality of the smallest chromatic class of $\varsigma$, that is, let $x=min\{\left|c^{-1}(i)\right|:i\in\left[k\right]\}$. Without loss of generality, suppose that $x=\left|c^{-1}(k)\right|$.
On the one hand, $c$ defines a partition of $E(K_{n})$ and it follows that $k\leq f_{n}(x):=n(n-1)/(2x)$.
On the other hand, $c$ is connected, therefore, $c^{-1}(k)$ is at least a tree subgraph. The maximum number of spanning trees in the subgraph $H=K_{x+1}$, such that $V(H)=V(c^{-1}(k))$, is at most $\frac{\binom{x+1}{2}}{x}$, then the chromatic class $x$ is incident to at most $\frac{\binom{x+1}{2}}{x}-1=\frac{x-1}{2}$ other chromatic classes. In addition, there are $(x+1)(n-(x+1))$ edges that are incident to some vertex of $c^{-1}(k)$ at least once, but on average each class is $\frac{2x}{x+1}$ incident to $c^{-1}(k)$. Since $c$ is complete, the number of chromatic classes incident to $c^{-1}(k)$ containing no edge in $H$ is at most
$\frac{(x+1)(n-x-1)}{2x/(x+1)}$ (6.1)
in average.
Hence, there are at most $g_{n}(x)-1$ chromatic classes incident with some edge in $c^{-1}(k)$ where $g_{n}(x)-1:=\frac{(x+1)^{2}}{2x}(n-(x+1))+\frac{x-1}{2}$, according to the hypothesis of the average degree, i.e.
$g_{n}(x)=\frac{n(x^{2}+2x+1)-x^{3}-2x^{2}-2x-1}{2x}.$
Then, we have $\psi_{c}(n)\leq k\approx\min\{f_{n}(x),g_{n}(x)\}$ and then
$\psi_{c}(n)\leq k\approx\max\left\{\min\{f_{n}(x),g_{n}(x)\}\textrm{ with }x\in\mathbb{N}\right\}.$
As $f_{n}$ is a hyperbola and $g_{n}$ is a parabola there are two positive values $x_{0}$ and $x_{1}$ such that $x_{0}<x_{1}$, $g_{n}(x_{0})=f_{n}(x_{0})$ and $g_{n}(x_{1})=f_{n}(x_{1})$ for which $f_{n}(x)\leq g_{n}(x)$ for $x_{0}\leq x\leq x_{1}$ and $g_{n}(x)<f_{n}(x)$ in any other case when $x\in\mathbb{R}^{+}$. Since $f_{n}(x_{0})>f_{n}(x_{1})$, it follows that $g_{n}(x_{0})=f_{n}(x_{0})=\max\left\{\min\{f_{n}(x),g_{n}(x)\}\textrm{ with }x\in\mathbb{R}^{+}\right\}$, see Figure 3.
Figure 3: The functions $g_{n}$ and $f_{n}$ for some fixed value $n$.
Equation $f_{n}(x_{0})=g_{n}(x_{0})$ is reduced to $n^{2}-(x_{0}^{2}+2x_{0}+2)n+x_{0}^{3}+2x_{0}^{2}+2x_{0}+1=0$. For the positive solution, we obtain $n=x_{0}^{2}+x_{0}+1$ and then $x_{0}=\frac{\sqrt{4n-3}-1}{2}$ taking the positive solution.
Since $f_{n}(x_{0})=\frac{n(n-1)}{2x_{0}}$ we get $f_{n}(x_{0})=\frac{n(n-1)}{\sqrt{4n-3}-1},$ therefore
$\psi_{c}(n)\leq k\approx\frac{n(n-1)}{\sqrt{4n-3}-1}$
and the first result is established.
For the second result, the hypothesis that each class contains at least one cycle implies that the average degree is at least 2. Equation 6.1 is modified as follows:
$\frac{(x+1)(n-x-1)}{2}$ (6.2)
Hence,
$g_{n}(x)=\frac{(n-x)(x+1)}{2}.$
In this case, equation $f_{n}(x_{0})=g_{n}(x_{0})$ is reduced to $n^{2}-n-x_{0}^{2}n-x_{0}n+x_{0}^{3}+x_{0}^{2}=0$. For the positive solution, we obtain $n=x_{0}^{2}$ (since $x_{0}\geq 3$, then $n\geq 9$) and then $x_{0}=\sqrt{n}$ taking the positive solution.
Since $f_{n}(x_{0})=\frac{n(n-1)}{2x_{0}}$, we get $f_{n}(x_{0})=\frac{n(n-1)}{2\sqrt{n}},$ and the last result is established. ∎
[…]
Figure 4: Improved bounds for $\psi_{c}(n)$.
$n$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$ Upper $1$ $3$ $4$ $6$ $7$ $10$ $14$ $18$ $22$ $25$ Approx Upper $12$ $15$ $17$ $20$ Lower Improved $13$ $16$ $18$ Upper Heuristic $12$ $14$ $16$ Lower $1$ $3$ $4$ $6$ $7$ $10$ $11$ $12$ $13$ $14$ $n$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$ $21$ Upper $28$ $31$ $34$ $37$ $40$ $45$ $51$ $57$ $63$ $70$ Approx Upper $23$ $26$ $28$ $32$ $35$ $38$ $41$ $45$ $48$ $52$ Lower Improved $21$ $32$ $35$ $37$ $39$ $43$ Upper Heuristic $19$ $21$ $24$ $27$ $30$ $32$ $36$ $39$ $42$ $45$ Lower $19$ $26$ $27$ $28$ $29$ $30$ $31$ $32$ $33$ $42$ $n$ $22$ $23$ $24$ $25$ $26$ $27$ $28$ $29$ $30$ $31$ Upper $74$ $78$ $82$ $86$ $90$ $94$ $98$ $102$ $106$ $116$ Approx Upper $56$ $59$ $63$ $67$ $71$ $75$ $80$ $84$ $88$ $93$ Lower Improved $46$ $50$ $52$ $54$ $55$ $59$ $62$ $64$ $66$ Upper Heuristic $49$ $52$ $56$ $60$ $63$ $67$ $71$ $75$ $79$ $83$ Lower $43$ $44$ $45$ $46$ $47$ $48$ $49$ $50$ $51$ $93$
Table 4: Improved bounds for $\psi_{c}(n)$.
As we can see in Table 4, for the values of $n=13$ and $31$ there are good approximations, which is due to the existence of the finite projective plane of odd order $q=3$ and $5$, which gives a lower bound of $\psi_{c}(n)$ for $n=q^{2}+q+1$ when $q$ is an odd prime power according to the following result.
Theorem 6.2.
[2] If $q$ is an odd prime power and $n=q^{2}+q+1$ then
$\left\lceil\frac{q}{2}\right\rceil n\leq\psi(n).$
Our approximation $\frac{n(n-1)}{\sqrt{4n-3}-1}$ implies that $\left\lceil\frac{q}{2}\right\rceil n$ is a good approximation because $q$ is odd and then
$\frac{n(n-1)}{\sqrt{4n-3}-1}=\frac{n(q^{2}+q)}{\sqrt{4(q^{2}+q+1)-3}-1}=\frac{q(q+1)n}{\sqrt{4q^{2}+4q+1}-1}=\frac{q(q+1)n}{2q}=\frac{(q+1)}{2}n=\left\lceil\frac{q}{2}\right\rceil n.$
7 Conclusions
The problem at hand involves improving, by adapting a genetic algorithm called Rank GA, the known bounds of the connected pseudoachromatic index, denoted $\psi_{c}^{\prime}(K_{n})$, which represents the maximum number of colors in a connected and complete coloring of a complete graph $K_{n}$ of order $n$. The Rank GA demonstrated significant effectiveness in improving $\psi_{c}^{\prime}(K_{n})$ bounds for complete graphs with connected classes. The algorithm was able to find the solutions due to its balance in exploring the solution space and exploiting promising solutions. The improved bounds of $\psi_{c}^{\prime}(K_{n})$ obtained by the Rank GA were validated against an analytical approximation, confirming their validity. The success of the Rank GA in enhancing the bounds of $\psi_{c}^{\prime}(K_{n})$ may significantly contribute to theoretical advancements in Chromatic Graph Theory. Upon examination of the resulting colorings, it is observed that the sizes of the chromatic classes vary, with most sizes equal to the minimum size stipulated by Theorem 6.1. However, their distribution does not align with the color patterns described in Theorem 6.2, meaning that they lack a projective plane structure in their arrangement.
Beyond the specific problem of graph coloring, the findings underscore the interdisciplinary impact of utilizing genetic algorithms in theoretical mathematical research.
Appendix A Example Solutions
Figure 5 presents a sample solution with $n=12$ and 21 colors. Additional examples and a Python-based visualizer can be downloaded from ***https://www.dropbox.com/scl/fo/ibudsyzgwj0gz8gzs8ifi/AEXSShINNc23KMsURaxS9w4?rlkey=jsgye9loui8gapqnbm89c0csq&dl=0. The visualizer allows users to toggle the visualization of each color class to analyze solutions.
Figure 5: Example Solution